In a t a=5 b=7 and sinA =3/4 then the number  of possible triangles are
 

Grade:11

1 Answers

Sai Soumya
5 years ago
dear student , 
it is given that a=5 , b=7 and sin A =3/4
as we know that , sin A / a=sin B/b
                       =>3/4/5=sin B/7
on solving we will get sin B=21/2o , which is not possible because its value if more than 1